# Simulation-based population inference of LISA’s Galactic binaries: Bypassing the global fit

### v1

@@ Abstract
The Laser Interferometer Space Antenna (LISA) is expected to detect thousands of individually resolved gravitational wave sources, overlapping in time and frequency, on top of unresolved astrophysical and/or primordial backgrounds. Disentangling resolved sources from backgrounds and extracting their parameters in a computationally intensive “global fit” is normally regarded as a necessary step toward reconstructing the properties of the underlying astrophysical populations. Here, we show that it is possible to infer the properties of the most numerous population of LISA sources – Galactic double white dwarfs – directly from the frequency (or, equivalently, time) strain series, by using a simulation-based approach that bypasses the global fit entirely. By training a normalizing flow on a custom-designed compression of simulated LISA frequency series from the Galactic double white dwarf population, we demonstrate how to infer the posterior distribution of population parameters (e.g., mass function, frequency, and spatial distributions). This allows for extracting information on the population parameters from both resolved and unresolved sources simultaneously and in a computationally efficient manner. Our approach to target population properties directly can be readily extended to other source classes (e.g., massive and stellar-mass black holes, extreme mass ratio inspirals), provided fast simulations are available, and to scenarios involving non-Gaussian or non-stationary noise (e.g., data gaps).
@@ I Introduction
The launch of the Laser Interferometer Space Antenna (LISA) LISA:2017pwj ; LISA:2024hlh , scheduled for the mid-2030s and led by ESA, in collaboration with NASA, will usher in a new era of precision gravitational-wave (GW) astronomy. Operating in the millihertz frequency range—well below that accessible to ground-based detectors such as LIGO, Virgo, and KAGRA—LISA will be sensitive to a wide variety of sources. These include massive black hole binaries (MBHBs) Klein:2015hvg ; Barausse:2023yrx , extreme mass ratio inspirals (EMRIs) Babak:2017tow , stellar-mass black hole binaries (stBHBs) Sesana:2016ljz , and an exceptionally large population of double white dwarfs (DWDs) Korol:2017qcx ; Korol+2022 ; Lamberts:2019nyk .
@@
In contrast to the transient signals detected by terrestrial observatories, many of LISA’s sources will be persistent, remaining in-band throughout the mission’s multi-year duration and overlapping in both time and frequency. The resulting data stream will be signal-dominated, with certain frequencies affected by a confusion foreground produced by millions of DWDs, mainly from our own Galaxy Korol:2017qcx ; Korol+2022 ; Lamberts:2019nyk . While this richness of signals carries a huge potential for deepening our understanding of the astrophysics of compact binaries, it also poses significant data analysis challenges. Disentangling and characterizing thousands of individually resolvable sources requires simultaneously fitting for an unknown number of overlapping sources – a task known as ‘the global fit’ Littenberg:2020bxy ; Lackeos:2023eub ; Littenberg:2023xpl ; Deng:2025wgk ; Strub:2024kbe ; Katz:2024oqg .
@@
Due to their large number, Galactic DWDs play a crucial role in the LISA global fit program. Not only do they provide a stochastic foreground, but several thousands of them are also expected to be individually resolvable Korol:2017qcx ; Korol+2022 ; Lamberts:2019nyk . From the astrophysical point of view, DWDs are relics of the binary stellar evolution process. Their GW signals encode valuable information about the end stages of stellar life, common envelope evolution, and the structure of the Milky Way LISA:2022yao . Understanding the complete population of DWDs, from resolved to unresolved systems, is therefore central to LISA’s scientific program.
@@
The traditional approach to astrophysical inference with LISA is based on the success of the global fit Littenberg:2020bxy ; Lackeos:2023eub ; Littenberg:2023xpl ; Deng:2025wgk ; Strub:2024kbe ; Katz:2024oqg which aims to simultaneously model all resolvable signals and estimate their parameters. In doing so, population inference biases must be avoided by accounting for the selection effects (Malmquist bias) that arise when fitting loud resolvable events and simultaneously modeling the unresolvable background (Adams+2012, ). Proposed global-fit methods are computationally intensive, involving thousands of parameters and potentially large degeneracies between them. In practice, one often relies on simplifying assumptions and iterative pipelines, whose scalability and accuracy remain a key concern. In the context of population inference, a faster, scalable estimation from simulation-based inference (SBI) SBIfrontier:Kranmer+2020 can complement the rigorous, yet slower, global fit.
@@
The global fit pipeline is typically optimized for source-by-source detection and cataloging Johnson:2025oyu . Although this enables detailed follow-up of high signal-to-noise sources, it introduces limitations for population-level inference. In a traditional framework, the parameters of the detected sources are first inferred and used to constrain population parameters, such as those that describe the mass function, frequency distribution or spatial density Korol_Rossi_Barausse_2018 (see also Adams:2012qw ; Delfavero_Breivik+2025 ). This two-step approach is vulnerable to selection biases, particularly in confusion-dominated regimes, and discards potentially valuable information from sources that fall below the detection threshold (i.e., from the foreground, which also contains information about population parameters Breivik+2020 ). Alternatively, performing both source and population inference simultaneously can be computationally prohibitive.
@@
In this work, we propose a fundamentally different strategy: to bypass source-level reconstruction altogether and perform population inference directly from the LISA time (or frequency) series. Our method is based on SBI SBIfrontier:Kranmer+2020 —also known as likelihood-free inference—which has emerged in recent years as a powerful alternative to traditional methods in scenarios where the likelihood is intractable or expensive to compute, but forward simulations are available. We apply SBI to the problem of Galactic DWDs, aiming to extract astrophysical population parameters directly from the raw LISA data, without characterizing individual sources. This differs from other neural network-based hierarchical inference methods, which infer population parameters from resolvable GW sources Leyde+2024 ; Mold+2025 .
@@
Our approach involves four components. First, using the prescription of Korol+2022 ; Korol+2022_code , we generate forward simulations of the Galactic DWD population using parameterized astrophysical prescriptions for the population’s frequency distribution, mass function, and spatial density within the Milky Way. We generate GWs, accounting for the LISA response, using the GBGPU implementation of the FastGB waveforms (Katz+2022, ; katz2022gbgpu, ; Cornish_Littenberg_2007, ) and the instrumental noise from the LISA mission requirement document (as implemented in the LDC software LDC2022 ) to produce realistic synthetic datasets. Second, we compress the data, extracting key features necessary for inference. Third, we train a masked-auto-regressive normalizing flow MAF —a deep neural network architecture that can model complex, high-dimensional probability distributions—to approximate the posterior over population parameters given observed data. Finally, we evaluate the quality of this inference, and apply posterior recalibration Crisostomi:2023tle techniques to ensure unbiased posterior estimates.
@@
Since simulation-based inference techniques can overfit or be miscalibrated, especially when faced with model degeneracies or limited training data (SNLE_2018, ; Talts+2018, ), we assess the statistical coverage of our posteriors for the population parameters across representative test sets, verifying whether true (simulated) parameters lie within the posterior intervals at the nominal credible rate. Where needed, we apply (approximate) Bayesian recalibration Crisostomi:2023tle to correct for under- or over-coverage, thus ensuring improved coverage.
@@
A key advantage of our method is that it exploits information from both resolved and unresolved sources simultaneously, making use of the full signal content. Moreover, once trained on simulated data, the method is amortized, meaning it can produce immediate posterior estimates for multiple datasets without any retraining or generation of extra simulations. The nearly instantaneous inference dramatically reduces the computational burden when LISA data are continuously received during the mission. Importantly, the method is also modular and extensible: it can be easily adapted to incorporate other LISA source classes–including MBHBs, stBHBs and EMRIs– and more realistic detector conditions involving non-Gaussian noise, data gaps, or time-dependent systematics Dey:2021dem ; Baghi:2019eqo ; Spadaro:2023muy ; Houba:2024ysm ; Burke:2025bun . While our present analysis focuses on Galactic binaries, the framework we propose lays the foundation for a unified treatment of all types of sources detectable by LISA.
@@ II Method
Our aim is to develop an amortized SBI framework to enable rapid posterior evaluation over a wide range of injected population parameter values. Once LISA starts its observation and its real data is available, non-amortized inference strategies, such as sequential SBI (SBIfrontier:Kranmer+2020, ; TMNRE_Miller+2021, ), can yield improved posterior accuracy due to optimized allocation of simulation resources around the observed data. Until real-data observations are available, the amortized approach facilitates rapid systematic benchmarking of inference performance and network optimization for a variety of potential observational data.
@@
Figure 1: Flowchart depicting the SBI training pipeline. The yellow arrows indicate the steps involved in training the deep neural network, and the blue arrows represent the calibration procedure. In the posterior plots, examples of the posterior probability of the injections $p_{\mathrm{SBI}}(\bm{\Lambda}_{\mathrm{GT}}|\bm{D})$ are shown by the green (red) circles representing cases where the injection lies within (outside) the contours of nominal coverage 0.8. The corresponding empirical coverage is marked in the calibration PP plot. The figure illustrating the simulated population is adapted from LSIA_redbook_2024 .
@@ II.1 Data Generation
As the name implies, SBI requires a dataset of simulations to train a deep neural network, with inference quality and fidelity dependent on the accuracy of the simulator. Therefore, developing a fast, reliable forward simulator that closely models observations is essential (Perigrine_2023, ). In the context of inferring the Galactic DWD population from LISA data, the forward simulator must first generate a plausible astrophysical population of the DWDs within the Milky Way, consistent with electromagnetic observations. Then, the simulator must produce the LISA data corresponding to the astrophysical sources and the expected instrumental noise. For this first attempt at DWD analysis, we ignore the presence of other sources (e.g. MBHBs, stBHBs, and EMRIs) within the LISA data.
@@ II.1.1 Population model
We adopt the Galactic DWD population simulator from the observation-driven model of Korol+2022 , which is based on the analysis of the spectroscopic samples from the Sloan Digital Sky Survey (SDSS) and the Supernova Ia Progenitor surveY (SPY) from Ref. Maoz+2018 . Interestingly, Korol+2022 finds that their estimates of the DWD population differ significantly from those predicted by binary population synthesis codes. The discrepancy between population synthesis estimates and observation-driven estimates, combined with the long runtime of the former, motivates us to build our forward simulator based on the latter model. We first summarize the parametric model below and later our GPU-optimized implementation.
Ref. Korol+2022 parameterizes the DWD population by their mass distribution, separation distribution, and their spatial number density in the Galaxy. The primary (more massive) star’s mass ($m_{1}$) distribution is modeled after the observed single white-dwarf mass distribution (Kepler+2015, ). Based on the observed flat mass ratio distribution Moe_DiStefano_2017 ; Duchene_Kraus_2013 , the secondary mass ($m_{2}$) is sampled from a uniform distribution in the range $[0.15\mathrm{M}_{\odot},m_{1}]$. Ref. Maoz+2018 finds that the distribution $n_{a}$ of the separation ($a$) of DWDs as a function of the time of coalescence $t_{0}$ and masses $m_{1}$, $m_{2}$ is well approximated by
@@
Eq. (1) is approximately a broken power-law with index $\alpha$ when the DWDs merger time is greater than the age of the Milky Way disk ($x\gg 1$). In the opposite limit $x\ll 1$, the power-law index is 3 for $\alpha\geq-1$, and $\alpha+4$ for $\alpha<-1$. To produce LISA data, the DWD separation is drawn from Eq. (1), with a minimum separation of 20,000 km (at merger) and a maximum corresponding to the LISA minimum frequency ($\sim 10^{-4}$ Hz).
@@
where $n_{0}=(4.49\pm 0.38)\times 10^{-3}\,\mathrm{pc}^{-3}$ is the local white dwarf density from the estimates of Hollands+2018 from the observational data of Gaia2018 . $R$ and $z$ are Galactocentric cylindrical coordinates, while $R_{\mathrm{d}}$ and $z_{\mathrm{d}}$ are free parameters describing the Milky Way disk’s scale width and height. To estimate the distance to the LISA detector, which enters in the waveform, one also needs the position of the Sun in Galactocentric coordinates, ($\mathrm{R}_{\odot}$, $\mathrm{z}_{\odot}$), which we fix to (8.1, 0.03) kpc Gravity:2019 .
Besides this disk population, we also consider a Galactic bulge population of DWDs (absent in Ref. Korol+2022 ). We follow Ref. Korol_Rossi_Barausse_2018 and parametrize its density $n_{\mathrm{b}}$ as
@@
$m_{\mathrm{0}}$ Primary-mass peak  [$\mathrm{M}_{\odot}$] $\mathcal{U}(0.15,1.4)$ 0.649
$\log_{10}m_{\mathrm{\gamma}}$ Primary-mass log-half width half maximum $\mathcal{U}[-2.0,0.5]$ 2.686
@@
$\alpha$ DWD separation index $\mathcal{U}(-2.05,-0.55)$1.3
@@
We adopt a uniform prior over our population parameters $\{m_{\mathrm{0}},\,\log_{10}m_{\mathrm{\gamma}},\,R_{\mathrm{d}},\,z_{\mathrm{d}},\,r_{\mathrm{b}},\,f_{\mathrm{b}},\,\alpha,\,\log_{10}N\}$. We set astrophysically agnostic prior limits on all parameters, such that they contain at least the 5-$\sigma$ observational electromagnetic constraints. $m_{\mathrm{0}}$ and $\log_{10}m_{\mathrm{\gamma}}$, the primary-mass distribution parameters, have values ranging from [0.15, 1.4] $\mathrm{M}_{\odot}$ and [-2, -0.5] respectively. The Galactic parameters $R_{\mathrm{d}}$, $z_{\mathrm{d}}$, $r_{\mathrm{b}}$, $f_{\mathrm{b}}$ have ranges [1.0, 4.0] kpc, [0.1, 1.0] kpc, [0.1, 1.0] kpc, and [0.05, 0.6] respectively based on MWdisclength_2008 ; MWscaleheight_2017 ; MWscalebulge_2009 . $\alpha$ has range [-2.05, -0.55] (Maoz+2018, ) and $\log_{10}N$ varies from [6.3, 7.8] based on the 5-$\sigma$ uncertainties on $\rho_{0}$ (Hollands+2018, ) and $f_{\mathrm{DWD}}$ (Korol+2022, ). We also define a fiducial value based on those considered in Korol+2022 ; Korol_Rossi_Barausse_2018 . Table 1 summarizes the population parameters, their descriptions, prior distributions, and fiducial values.
For our analysis, we adopt the codes of Korol+2022_code ; natalia_code_2022 based on the model from Korol+2022 , but with our additional Milky Way bulge population and simplified parametric mass distribution. Our novel GPU-accelerated implementation [^1] Code to be released shortly. that leverages the computing facilities of CuPy and Numba, a just-in-time (JIT) compiler, yields three orders of magnitude faster population generation than Korol+2022_code . For a given set of parameter values, the code can produce a Galactic realization in about a minute. This enables us to simulate millions of possible DWD populations, each with tens of millions of DWDs.
Fig. 2 illustrates the frequency distribution of three realizations of the DWD population from different primary-mass distributions and other parameters fixed at a fiducial value. The inset shows the absolute value of the frequency spectrum generated for the three DWD populations, differing only in their primary mass distribution. We overplot the LISA typical instrumental noise amplitude spectral density (ASD) for reference. The vast number of overlapping sources at low frequencies ($10^{-4}-10^{-3}$ Hz) contribute to the Galactic “foreground” or “confusion noise”, an unfortunate term considering that some of the parameter inference signature could potentially lie here. We explore this in further detail in § IV.

### v2

# Simulation-based population inference of LISA's Galactic binaries: Bypassing the global fit

@@ Abstract
The Laser Interferometer Space Antenna (LISA) is expected to detect thousands of individually resolved gravitational wave sources, overlapping in time and frequency, on top of unresolved astrophysical and/or primordial backgrounds. Disentangling resolved sources from backgrounds and extracting their parameters in a computationally intensive ``global fit'' is normally regarded as a necessary step toward reconstructing the properties of the underlying astrophysical populations. Here, we show that it is in principle feasible to infer the population properties of the most numerous of LISA sources – Galactic double white dwarfs – directly from the frequency (or, equivalently, time) strain series by adopting a simulation-based approach, without extracting and estimating the parameters of each single source. By training a normalizing flow on a custom-designed compression of simulated LISA frequency series from the Galactic double white dwarf population, we demonstrate how to infer the posterior distribution of population parameters (e.g., mass function, frequency, and spatial distributions). This allows for extracting information on the population parameters from both resolved and unresolved sources simultaneously and in a computationally efficient manner. This approach can be extended to other source classes (e.g., massive and stellar-mass black holes, extreme mass ratio inspirals) and to scenarios involving non-Gaussian or non-stationary noise (e.g., data gaps), provided that fast and accurate simulations are available.
@@ I Introduction
The launch of the Laser Interferometer Space Antenna (LISA) Amaro-Seoane and others (2017); Colpi and others (2024), scheduled for the mid-2030s and led by ESA, in collaboration with NASA, will usher in a new era of precision gravitational-wave (GW) astronomy. Operating in the millihertz frequency range—well below that accessible to ground-based detectors such as LIGO, Virgo, and KAGRA—LISA will be sensitive to a wide variety of sources. These include massive black hole binaries (MBHBs) Klein and others (2016); Barausse et al. (2023), extreme mass ratio inspirals (EMRIs) Babak et al. (2017), stellar-mass black hole binaries (stBHBs) Sesana (2016), and an exceptionally large population of double white dwarfs (DWDs) Korol et al. (2017, 2022c); Lamberts et al. (2019).
@@
In contrast to the transient signals detected by terrestrial observatories, many of LISA's sources will be persistent, remaining in-band throughout the mission's multi-year duration and overlapping in both time and frequency. The resulting data stream will be signal-dominated, with certain frequencies affected by a confusion foreground produced by millions of DWDs, mainly from our own Galaxy Korol et al. (2017, 2022c); Lamberts et al. (2019). While this richness of signals carries a huge potential for deepening our understanding of the astrophysics of compact binaries, it also poses significant data analysis challenges. Disentangling and characterizing thousands of individually resolvable sources requires simultaneously fitting for an unknown number of overlapping sources – a task known as `the global fit' Littenberg et al. (2020); Lackeos et al. (2023); Littenberg and Cornish (2023); Deng et al. (2025); Strub et al. (2024); Katz et al. (2025).
@@
Due to their large number, Galactic DWDs play a crucial role in the LISA global fit program. Not only do they provide a stochastic foreground, but several thousands of them are also expected to be individually resolvable Korol et al. (2017, 2022c); Lamberts et al. (2019). From the astrophysical point of view, DWDs are relics of the binary stellar evolution process. Their GW signals encode valuable information about the end stages of stellar life, common envelope evolution, and the structure of the Milky Way Seoane and others (2023). Understanding the complete population of DWDs, from resolved to unresolved systems, is therefore central to LISA's scientific program.
@@
The traditional approach to astrophysical inference with LISA is based on the success of the global fit Littenberg et al. (2020); Lackeos et al. (2023); Littenberg and Cornish (2023); Deng et al. (2025); Strub et al. (2024); Katz et al. (2025) which aims to simultaneously model all resolvable signals and estimate their parameters. In doing so, population inference biases must be avoided by accounting for the selection effects (Malmquist bias) that arise when fitting loud resolvable events and simultaneously modeling the unresolvable background (Adams et al., 2012b). Proposed global-fit methods are computationally intensive, involving thousands of parameters and potentially large degeneracies between them. In practice, one often relies on simplifying assumptions and iterative pipelines, whose scalability and accuracy remain a key concern. In the context of population inference, a global fit procedure can only infer population properties after estimating the parameters of the resolvable sources and the foreground. In contrast, a faster, scalable approach based on simulation-based inference (SBI) Cranmer et al. (2020) can directly estimate population properties without the need for source-level inference.
@@
The global fit pipeline is typically optimized for source-by-source detection and cataloging Johnson et al. (2025). Although this enables detailed follow-up of high signal-to-noise sources, it introduces limitations for population-level inference. In a traditional framework, the parameters of the detected sources are first inferred and used to constrain population parameters, such as those that describe the mass function, frequency distribution, or spatial density Korol et al. (2019) (see also Adams et al. (2012a); Delfavero et al. (2025)). This two-step approach is vulnerable to selection biases, particularly in confusion-dominated regimes, and discards potentially valuable information from sources that fall below the detection threshold (i.e., from the foreground, which also contains information about population parameters Breivik et al. (2020)). Alternatively, performing both source and population inference simultaneously within the global fit can be computationally expensive. In contrast, our method estimates the population properties without the need for source-level inference.
@@
In this work, we propose a fundamentally different strategy: to bypass source-level reconstruction altogether and perform population inference directly from the LISA time (or frequency) series. Our method is based on SBI Cranmer et al. (2020)—also known as likelihood-free inference—which has emerged in recent years as a powerful alternative to traditional methods in scenarios where the likelihood is intractable or expensive to compute, but forward simulations are available. We apply SBI to the problem of Galactic DWDs, aiming to extract astrophysical population parameters directly from the raw LISA data, without characterizing individual sources. This differs from other neural network-based hierarchical inference methods, which infer population parameters from resolvable GW sources Leyde et al. (2024); Mould et al. (2025).
@@
Our approach involves four components. First, using the prescription of Korol et al. (2022c, b), we generate forward simulations of the Galactic DWD population using parameterized astrophysical prescriptions for the population's frequency distribution, mass function, and spatial density within the Milky Way. We generate GWs, accounting for the LISA response, using the GBGPU implementation of the FastGB waveforms (Katz et al., 2022; Katz, 2022; Cornish and Littenberg, 2007a) and the instrumental noise from the LISA mission requirement document (as implemented in the LDC software LISA Data Challenge working group (2022)) to produce realistic synthetic datasets. Second, we compress the data, extracting key features necessary for inference. Third, we train a masked-auto-regressive normalizing flow Papamakarios et al. (2017)—a deep neural network architecture that can model complex, high-dimensional probability distributions—to approximate the posterior over population parameters given observed data. Finally, we evaluate the quality of this inference, and apply posterior recalibration Crisostomi et al. (2023) techniques to ensure unbiased posterior estimates.
@@
Since simulation-based inference techniques can overfit or be miscalibrated, especially when faced with model degeneracies or limited training data (Papamakarios et al., 2018; Talts et al., 2018), we assess the statistical coverage of our posteriors for the population parameters across representative test sets, verifying whether true (simulated) parameters lie within the posterior intervals at the nominal credible rate. Where needed, we apply (approximate) Bayesian recalibration Crisostomi et al. (2023) to correct for under- or over-coverage, thus ensuring improved coverage. These checks are conditional on modeling the data correctly. In the results presented in this paper, that assumption is automatically satisfied because the observations are generated using the same simulator as the training data, but more powerful tests will be required Schmitt et al. (2021); Huang et al. (2023a); Wehenkel et al. (2024) to deploy our method on actual LISA data.
@@
A key advantage of our method is that it exploits information from both resolved and unresolved sources simultaneously, making use of the full signal content (suitably compressed). Moreover, once trained on simulated data, the method is amortized, meaning it can produce immediate posterior estimates for multiple datasets without any retraining or generation of extra simulations. This enables astrophysical inference directly from the data stream—without relying on, or waiting for, global-fit results on individually resolved sources. Importantly, the method is extensible: it can be easily adapted to incorporate other LISA source classes–including MBHBs, stBHBs and EMRIs – and more realistic detector conditions involving non-Gaussian noise, data gaps, or time-dependent systematics Dey et al. (2021); Baghi et al. (2019); Spadaro et al. (2023); Houba et al. (2024); Burke et al. (2025). While our present analysis focuses on Galactic binaries and Gaussian noise for ease of simulation and the purpose of illustrating the approach, the framework we propose lays the foundation for a unified treatment of all types of sources detectable by LISA. For example, in future analyses that include multiple source classes, we can re-use pre-computed simulations of GWs from the DWD population. This ensures that the simulation budget does not scale dramatically, but only increases to the extent necessary to provide adequate coverage of the additional source populations. Several challenges, however, remain to be addressed before this can be achieved, and we discuss them in the Conclusions.
@@ II Method
Our aim is to develop an amortized SBI framework to enable rapid posterior evaluation over a wide range of injected population parameter values. Once LISA starts its observation and its real data is available, non-amortized inference strategies, such as sequential SBI (Cranmer et al., 2020; Miller et al., 2021), can yield improved posterior accuracy due to optimized allocation of simulation resources around the observed data. In this work, however, we limit ourselves to a faster, amortized, and non-sequential approach, in order to systematically assess the performance of our method across the population parameter space. An extension of our method to focus on a particular data set (i.e., when LISA begins observation) can be easily achieved with a multi-round inference approach.
@@
Figure 1: Flowchart depicting the SBI training pipeline. The yellow arrows indicate the steps involved in training the deep neural network, and the blue arrows represent the calibration procedure. In the posterior plots, examples of the posterior probability of the injections $p_{\mathrm{SBI}}(\bm{\Lambda}_{\mathrm{GT}}|\bm{D})$ are shown by the green (red) circles representing cases where the injection lies within (outside) the contours of nominal coverage 0.8. The corresponding empirical coverage is marked in the calibration $P$ plot. The figure illustrating the simulated population is adapted from Colpi et al. (2024).
@@ II.1 Data Generation
As the name implies, SBI requires a dataset of simulations to train a deep neural network, with inference quality and fidelity dependent on the accuracy of the simulator. Therefore, developing a fast, reliable forward simulator that closely models observations is essential (Bhardwaj et al., 2023). In the context of inferring the Galactic DWD population from LISA data, the forward simulator must first generate a plausible astrophysical population of the DWDs within the Milky Way, consistent with electromagnetic observations. Then, the simulator must produce the LISA data corresponding to the astrophysical sources and the expected instrumental noise. For this first attempt at DWD analysis, we ignore the presence of other sources (e.g. MBHBs, stBHBs, and EMRIs) within the LISA data.
@@ II.1.1 Population model
We adopt the Galactic DWD population simulator from the observation-driven model of Korol et al. (2022c), which is based on the analysis of the spectroscopic samples from the Sloan Digital Sky Survey (SDSS) and the Supernova Ia Progenitor surveY (SPY) from Ref. Maoz et al. (2018). Interestingly, Korol et al. (2022c) finds that their estimates of the DWD population differ significantly from those predicted by binary population synthesis codes. The discrepancy between population synthesis estimates and observation-driven estimates, combined with the long runtime of the former, motivates us to build our forward simulator based on the latter model. We first summarize the parametric model below and later our GPU-optimized implementation. Our astrophysical model for the DWD population is intentionally simplified (to allow for fast simulations). For instance, we model the primary-mass distribution with a two-parameter Lorentzian. This is suitable for a proof-of-concept, but if the true distribution exhibits structure not captured by this form (e.g., sharp gaps or multimodality), the resulting inference may be systematically biased. Extending to more flexible models is straightforward but increases dimensionality. In such cases, dimensionality-reduction techniques (e.g., PCA), as in Taylor and Gerosa (2018), can mitigate complexity while retaining the dominant modes of variation. Alternatively, performing simulations from astrophysical population synthesis models is conceptually possible, but requires a dramatic speed up of the latter (via e.g. emulators).
Ref. Korol et al. (2022c) parameterizes the DWD population by their mass distribution, separation distribution, and their spatial number density in the Galaxy. The primary (more massive) star's mass ($m_{1}$) distribution is modeled after the observed single white-dwarf mass distribution (Kepler et al., 2015). Based on the observed flat mass ratio distribution Moe and Di Stefano (2017); Duch^ene and Kraus (2013), the secondary mass ($m_{2}$) is sampled from a uniform distribution in the range $0.15\mathrm{M}_{\odot},m_{1}]$. Ref. Maoz et al. (2018) finds that the distribution $n_{a}$ of the separation ($a$) of DWDs as a function of the age of the Milky Way $t_{0}$ and masses $m_{1}$, $m_{2}$ is well approximated by
@@
Eq. (1) is approximately a broken power-law with index $\alpha$ when the DWDs merger time is greater than the age of the Milky Way disk ($x\gg 1$). In the opposite limit $x\ll 1$, the power-law index is 3 for $\alpha\geq-1$, and $\alpha+4$ for $\alpha<-1$. To produce LISA data, the DWD separation is drawn from Eq. (1), with a minimum separation of 20,000 km, and a maximum corresponding to the LISA minimum frequency ($\sim 10^{-4}$ Hz).
@@
where $n_{0}=(4.49\pm 0.38)\times 10^{-3}\,\mathrm{pc}^{-3}$ is the local white dwarf density from the estimates of Hollands et al. (2018) from the observational data of Gaia Collaboration et al. (2018). $R$ and $z$ are Galactocentric cylindrical coordinates, while $R_{\mathrm{d}}$ and $z_{\mathrm{d}}$ are free parameters describing the Milky Way disk's scale width and height. To estimate the distance to the LISA detector, which enters in the waveform, one also needs the position of the Sun in Galactocentric coordinates, ($\mathrm{R}_{\odot}$, $\mathrm{z}_{\odot}$), which we fix to (8.1, 0.03) kpc Abuter and others (2019).
Besides this disk population, we also consider a Galactic bulge population of DWDs (absent in Ref. Korol et al. (2022c)). We follow Ref. Korol et al. (2019) and parametrize its density $n_{\mathrm{b}}$ as
@@
$m_{\mathrm{0}}$ Primary-mass peak $\mathrm{M}_{\odot}$] $\mathcal{U}(0.15,1.4)$ 0.649
$\log_{10}m_{\mathrm{\gamma}}$ Primary-mass log half-width at half-maximum $\mathcal{U}(-2.0,0.5)$ 2.686
@@
$\alpha$ DWD separation index $\mathcal{U}(-2.05,-0.55)$ -1.3
@@
We adopt a uniform prior over our population parameters $\{m_{\mathrm{0}},\,\log_{10}m_{\mathrm{\gamma}},\,R_{\mathrm{d}},\,z_{\mathrm{d}},\,r_{\mathrm{b}},\,f_{\mathrm{b}},\,\alpha,\,\log_{10}N\}$. We set astrophysically agnostic prior limits on all parameters, such that they contain at least the 5-$\sigma$ observational electromagnetic constraints. $m_{\mathrm{0}}$ and $\log_{10}m_{\mathrm{\gamma}}$, the primary-mass distribution parameters, have values ranging from [0.15, 1.4] $\mathrm{M}_{\odot}$ and [-2, -0.5] respectively. The Galactic parameters $R_{\mathrm{d}}$, $z_{\mathrm{d}}$, $r_{\mathrm{b}}$, $f_{\mathrm{b}}$ have ranges [1.0, 4.0] kpc, [0.1, 1.0] kpc, [0.1, 1.0] kpc, and [0.05, 0.6] respectively based on Juric et al. (2008); Mackereth et al. (2017); Sofue et al. (2009). $\alpha$ has range [-2.05, -0.55] (Maoz et al., 2018) and $\log_{10}N$ varies from [6.3, 7.8] based on the 5-$\sigma$ uncertainties on $\rho_{0}$ (Hollands et al., 2018) and $f_{\mathrm{DWD}}$ (Korol et al., 2022c). We also define a fiducial value based on those considered in Korol et al. (2022c, 2019). Table 1 summarizes the population parameters, their descriptions, prior distributions, and fiducial values.
For our analysis, we adopt the codes of Korol et al. (2022b); Korsakova (2022) based on the model from Korol et al. (2022c), but with our additional Milky Way bulge population and simplified parametric mass distribution. Our novel GPU-accelerated implementation [^1] Code to be released shortly. that leverages the computing facilities of CuPy and Numba, a just-in-time (JIT) compiler, yields three orders of magnitude faster population generation than Korol et al. (2022b). For a given set of parameter values, the code can produce a Galactic realization in about a minute. This enables us to simulate millions of possible DWD populations, each with tens of millions of DWDs.
Fig. 2 illustrates the frequency distribution of three realizations of the DWD population from different primary-mass distributions and other parameters fixed at a fiducial value. The inset shows the absolute value of the frequency spectrum generated for the three DWD populations, differing only in their primary mass distribution. We overplot the LISA typical instrumental noise amplitude spectral density (ASD) for reference. The vast number of overlapping sources at low frequencies ($10^{-4}-10^{-3}$ Hz) contributes to the Galactic ``foreground'' or ``confusion noise'', an unfortunate term considering that some of the parameter inference signature could potentially lie here. We explore this in further detail in § IV.